OpenStudy (anonymous):

find g'(4) given that f(4)=3 and f'(4) = -5. g(x) = f(x)/x

OpenStudy (anonymous):

g'(x) =( x f'(x) - f(x) ) / x^2 g'(4) = {4 * -5 - 3 ) / 4^2 g'(4) = -23/16